OpenStudy (anonymous):

Jaycee follows baseball closely. This season, his favorite professional player has a batting average of 0.250 or 25%. This means that, on average, he gets one hit every four times he is at bat. Jaycee wants to know the likelihood that his favorite player will get at least two hits in the five times he'll be at bat. So Jaycee conducted a simulation to find out.

OpenStudy (anonymous):

Its certain tht he wld have hit a ball in d first four balls .. for him to have atleast 2 shots his first ball should be hit in next four which has d probability of 1/4

OpenStudy (anonymous):

\[P(X=x)=nCr(5,x) p^x(1-p)^{5-x}\] where p is probability of getting a hit, and x is amount of hits out of 5 trials so you need to take 1-((P(X=0)+P(x=1))

OpenStudy (anonymous):

its not certain that he would have hit a ball in the first four or 5 or even 10000 at bats technically

OpenStudy (anonymous):

I get .367188 as the chance he will get a hit 2 or more of the 5 tries

OpenStudy (anonymous):

No its given in q tat he s sure to hit a ball in four balls ... so its d matter of nxt four balls and to get two hits in fours .. he should hit d first ball o d next four

OpenStudy (anonymous):

it says on average he gets one hit every 4 at bats, the question you posted never said he will surely get one hit every 4 times. You need to use a binomial distribution